1. Given the equation of the ellipse (1 – 4)2/25 + (y + 2)2/49

  2. a. The Center

  3. b. The Foci

  4. c. Width of major axis
Answer #1

(1-4)² t (7+2)2 49 25 a) centre =(4,-2) (6) foci cz √49-25 N24 i foci are (4, -24/24), (1, -2-v74) Major axisa 2x7 = 14